Exercise 10-2 Direct Labor Variances [LO10-2] SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 7,200 of these meals using 3,500 direct labor-hours. The company paid its direct labor workers a total of $31,500 for this work, or $9.00 per hour. According to the standard cost card for this meal, it should require 0.50 direct labor-hours at a cost of $8.50 per hour.

Required: 1. What is the standard labor-hours allowed (SH) to prepare 7,200 meals? 2. What is the standard labor cost allowed (SH × SR) to prepare 7,200 meals? 3. What is the labor spending variance? 4. What is the labor rate variance and the labor efficiency variance?
Answer #1

Standard hours allowed = 7200*0.50

= 3600

Standard cost allowed = 3600*8.50

= 30,600

Labor spending variance = 30,600 - 31,500

= 900 U

Labor Rate Variance = (SR-AR) *AH

= (8.50 - 9)*3500

= 1750 U

Labor efficiency variance = (SH-AH) *SR

= (3600-3500)*8.50

= 850 F